Show that the sequence (an) is bounded iff |an| is bounded .


1
Expert's answer
2022-04-12T12:14:25-0400

ANSWER

A sequence "\left \{a _{n} \right \}" is called bounded if there are real numbers "l" and "b" , such that

"l\leq a_{n}\leq b" for all "n\in N" . (1)

1) Let "\left \{a _{n} \right \}" be a bounded sequence, denote "M=max\left \{ |l |, \: |b|\right \}"

Since "|b|\leq M, |l|\leq M" , then "-M\leq l\leq M, -M\leq b\leq M" . Hence

"-M\leq l\leq a_{n}\leq b\leq M" for all "n\in N" .

It means, that

"0\leq |a_{n}|\leq M" (2)


Those sequence "\left \{|a _{n}| \right \}" is bounded.

2) Conversely , let the sequence "\left \{|a _{n}| \right \}" is bounded. Then exists "M>0" such that inequality (2)

holds for all "n\in N" ..Since (2) is equivalent to the inequality

"-M\leq a_{n}\leq M" ,

then (1) is satisfied for "l=-M, b=M" . Hence, the sequence "\left \{a _{n} \right \}" is bounded.